A car moving at 40 km/h is to be stopped applying brakes in the next 4.0 m. If the car weighs 2000 kg what average force must be applied on it?

Grade:11

1 Answers

Mustafa Ahmed
7 years ago
30864.2 N is correct?If yes then see how:Use newtons 2nd eqn of motionV^2 = u^2 - 2as(U = 40kph = 11.11 mps)Since v=0 hence a=-15.432mpssNow, average force = mass × acceln= 20000 × 15.432 = 30864.2